A charge of -7.4 µC is traveling at a speed of 8.6 106 m/s in a region of space where there is a magnetic field. The angle between the velocity of the charge and the field is 52°. A force of magnitude 4.7 10-3 N acts on the charge. What is the magnitude of the magnetic field?
Answer #1

Force on the electron in the magnetic field

F = Q*v*B *sin(\theta )

Q - charge = 7.4 *10^{-6}

v - velocity = 8.6x10^6

F- Force = 4.7x10^{-3} N

B- Mgnetic Filed = ??

Theta = 52^{\circ}

So B = \frac{F}{v*Q*sin(52^{\circ})}

B = \frac{4.7x10^{-3} N}{(8.6x10^6)(7.4 *10^{-6})sin(52^{\circ})}


B = 9.372 x 10^{-5} T